Find the equation for an exponential function that passes through the pair of points given below.
Through:(−2, 5) and (3, 3)
F(X)=

Answer by xinxin(76)

The basic expression for exponential function is F(x) = aK^x, plug in the numbers:
aK^(-2)=5
aK^3 = 3
From the first equation, substitute a with K: a = 5/K^(-2). Combine this with the second equation:5/K^(-2)*K^3 =3
Solve for K, get K = 5th root of 3/5 = 0.9, a = 4.1